OpenStudy (anonymous):

Simplify 7 log3 k + 6 log3 m – 9 log3 n? can someone show the steps on how you do this?

OpenStudy (anonymous):

A. log3 B. 4 log3 C. log3 D. 4 log3 (k + m – n)

OpenStudy (anonymous):

Let's first recall a couple of log identities: \[\left[\begin{matrix}\text{First} & x*\log(y) = \log(y^x) \\ \text{Second }& \log(x*y) = \log(x) + \log(y) \\\text{Third} & \log({x \over y}) = \log(x) - \log(y) \end{matrix}\right]\]From the First identity:\[\log(k^7) + \log(m^6) - \log(n^9)\]From the Second and Third Identity\[\log({k^7 m^6 \over n^9})\]

OpenStudy (anonymous):

hey eashmore, could you look at problem D on greenday's A 1100 kg car is traveling at 24 m/s. post?

OpenStudy (anonymous):

that is choice C, but it has 3with the log?

OpenStudy (anonymous):

Same difference. I omitted the base for ease of typing it up. Remember these identities only apply when the base is the same.

OpenStudy (anonymous):

so are you master of the logarithms now?

OpenStudy (anonymous):

not totally..lol

OpenStudy (anonymous):

keep it up.. have fun

OpenStudy (anonymous):

lol..do you know a good website to graph functions?

OpenStudy (anonymous):

http://www.wolframalpha.com

OpenStudy (anonymous):

What is the function? I'll type it in so you have an idea of how to type things in.

OpenStudy (anonymous):

oh yeah...totally forgot about tht..thanks

OpenStudy (anonymous):

thanks eashmore for the physics help

OpenStudy (anonymous):

f(x) = log0.3 x

OpenStudy (anonymous):

A base of 0.3?

OpenStudy (anonymous):

yes